OpenStudy (anonymous):

find the perimeter of a triangle with vertices (2,4),(1,3),and (2,-3)

OpenStudy (anonymous):

A(2,4) and B(1,3) x2=1 x1=2 y2=3 y1=4 So, (1-2)^2 + (3-4)^2 = 1+1=2 square root B(1,3) and C(2,-3) so, (2-1)^2 + (-3-3)^2 = 1+36=37 square root. A(2,4) and C(2,-3) (2-2)^2 + (-3-4)^2 = 0 + 49 = square root of 49 is 7 so Sides are \[7, \sqrt{37}, \sqrt{2}\] For the perimeter, just add them: \[\sqrt{37}+\sqrt{2}+7\] units.

OpenStudy (anonymous):

why do you do the square roots?

OpenStudy (anonymous):

\[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] This is the formula for finding the length of a side, so: \[\sqrt{(1-2)^{2}+(3-4)^{2}}\] \[\sqrt{1+1}=\sqrt{2}\] That's how we get the square roots :)

OpenStudy (anonymous):

ohh okay thank you!

OpenStudy (anonymous):

My pleasure! :)